# DMSC: Dynamic Multi-Scale Coordination Framework for Time Series Forecasting

### v2

@@ Abstract
Time Series Forecasting (TSF) faces persistent challenges in modeling intricate temporal dependencies across different scales. Despite recent advances leveraging different decomposition operations and novel architectures based on CNN, MLP or Transformer, existing methods still struggle with static decomposition strategies, fragmented dependency modeling, and inflexible fusion mechanisms, limiting their ability to model intricate temporal dependencies. To explicitly solve the mentioned three problems respectively, we propose a novel Dynamic Multi-Scale Coordination Framework (DMSC) with Multi-Scale Patch Decomposition block (EMPD), Triad Interaction Block (TIB) and Adaptive Scale Routing MoE block (ASR-MoE). Specifically, EMPD is designed as a built-in component to dynamically segment sequences into hierarchical patches with exponentially scaled granularities, eliminating predefined scale constraints through input-adaptive patch adjustment. TIB then jointly models intra-patch, inter-patch, and cross-variable dependencies within each layer’s decomposed representations. EMPD and TIB are jointly integrated into layers forming a multi-layer progressive cascade architecture, where coarse-grained representations from earlier layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. And ASR-MoE dynamically fuses multi-scale predictions by leveraging specialized global and local experts with temporal-aware weighting. Comprehensive experiments on thirteen real-world benchmarks demonstrate that DMSC consistently maintains state-of-the-art (SOTA) performance and superior computational efficiency for TSF tasks. Code is available at https://github.com/1327679995/DMSC.
@@ Introduction
Time Series Forecasting (TSF) constitutes a pivotal capability across numerous fields such as energy consumption(Alvarez et al. 2010; Guo et al. 2015), healthcare monitoring(Tran, Nguyen, and Shahabi 2019; Wei et al. 2022), transportation scheduling(Guo et al. 2020; Jin et al. 2021), weather forecasting(Wu et al. 2023b; Bi et al. 2023), economics(Chen et al. 2023b; Yu et al. 2023) and urban computing(Li et al. 2024; Yuan et al. 2024). Time series data exhibit intricate temporal dependencies, encompassing nonlinear relationships, long-term and short-term patterns such as trends, seasonality and dynamic variable couplings(Qiu et al. 2024). Owing to non-stationarity(Nason 2006) and complex interconnections(Dagum and Bianconcini 2016), traditional approaches struggle to effectively deal with time series data, making effectively dependency modeling a critical challenge for enhancing predictive precision(Shao et al. 2025). In recent years, the powerful ability of deep learning to capture complex dependency relationships in time series data has been demonstrated, making it a formidable tool in the field of TSF(Chen et al. 2023a).
@@
However, intricate dependencies in time series inherently manifest across multiscales: coarser granularities encapsulate long-term trends while finer resolutions capture short-term fluctuations and periodic patterns. Existing single-scale modeling approaches often fail to balance fine-grained local details and holistic global trends(Wang et al. 2025). To effectively model multiscale temporal dependencies, recent research has pioneered multiscale modeling frameworks that concurrently extract global and local patterns, thereby mitigating the representational limitations of single-scale approaches. While previous implementations often employ moving average downsampling(Wang et al. 2024b; Huang et al. 2025), which often remains insufficient in capturing local semantic information. In contrast, more effective alternatives focus on patch-wise and variable-wise decomposition operations that decompose sequence across feature and temporal dimensions. Such decomposition operations better preserves both global trends and local semantics, and multi-length patches inherently learn multi-scale representations. Critically, prevailing methods exhibit two constraints:1) dependence on fixed-scale decomposition strategies that lack dynamic scale adaptation; 2) fragmented processing of temporal dependencies and cross-variable interactions, undermining holistic dependency modeling.
@@
The core value of multiscale feature extraction resides in mining complementary information across diverse temporal granularities and hierarchical levels, thereby overcoming the limitations of single-scale modeling. Following feature capture, the effective design of prediction heads for multiscale fusion is equally important. Conventional approaches typically employ either linear projection layers(Li, Li, and Diao 2025), additive combinations of multiple projections(Wang et al. 2024b), or spectral amplitude-weighted aggregation(Wu et al. 2023a). However, above simplistic fusion method exhibit fundamental inadequacies: they fail to leverage heterogeneous dominance patterns across multiscales, disregard dynamic inter-scale importance, and induce quadratic parameter growth with increasing input lengths and scale counts, severely compromising deployment efficiency(Qiu et al. 2024).
@@
To effectively modeling complex temporal dependencies which inherently manifest across multiple patterns and scales, we propose a Dynamic Multi-Scale Coordination framework (DMSC) for TSF. This framework dynamically decomposes time series and extracts features across scales. It effectively captures intricate dependencies across temporal resolutions and cross-variable interactions. Meanwhile, it can adaptively predict learned representations through temporal-aware fusion. Specifically, DMSC processes intricate dependencies through a multi-layer progressive architecture and Adaptive Scale Routing Mixture-of-Experts (ASR-MoE). The multi-layer progressive cascade architecture incorporates Embedded Multi-Scale Patch Decomposition Block (EMPD), Triad Interaction Block (TIB), where EMPD transforms input into 3D representations at varying granularities and TIB models heterogeneous dependencies within each layer. In this architecture, coarse-grained representations from shallow layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. After multi-scale feature extraction, ASR-MoE is employed to perform time-aware weighted dynamic fusion for multi-scale prediction, enabling adaptive integration of features across scales. By synergistically integrating dynamic patch decomposition, deeper models and sparse MoE principles, DMSC achieves full-spectrum multi-scale coordination across embedding, feature extraction, and prediction stages. This dynamic co-optimization enable it to achieve SOTA performance across 13 real-world benchmarks while maintaining high efficiency and low computational complexity. Our contributions are as follows:
@@
  • A novel multi-layer progressive cascade architecture is designed to jointly integrate Embedded Multi-Scale Patch Decomposition Block (EMPD) and Triad Interaction Block (TIB) across layers. Unlike fixed-scale approaches, EMPD dynamically adjusts patch granularities by employing a lightweight neural controller based on input characteristics, and TIB jointly models intra-patch, inter-patch, and inter-variable dependencies through gated feature fusion to form a coarse-to-fine feature pyramid, where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated residual pathways.
@@
  • Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) is proposed to resolve static fusion limitations in prediction. ASR-MoE establishes a hierarchical expert architecture to explicitly decouples long-term and short-term dependencies, where global-shared experts capture common long-term temporal dependencies, while local-specialized experts model different short-term variations. Significantly, a temporal-aware weighting aggregator is designed to dynamically compute scale-specific prediction contributions with historical memory.
@@
Existing deep models for TSF tasks can be broadly categorized into MLP-based, CNN-based, and Transformer-based architectures(Wang et al. 2024c). MLP-based models(Oreshkin et al. 2020; Yi et al. 2023; Chi et al. 2021) primarily model temporal dependencies through predefined decomposition schemes and fully-connected architectures. While demonstrating simplicity and efficiency in capturing intra-series patterns, they exhibit critical limitations in modeling long-term dependencies and complex inter-series relationships. In contrast, CNN-based models(Bai, Kolter, and Koltun 2018; LIU et al. 2022; Wang et al. 2023) leverage different convolutions to effectively capture local dependencies. Although they remain constrained in modeling global relationships, recent efforts have mitigated this through large-kernel convolutions(Luo and Wang 2024) and frequency-domain analysis(Wu et al. 2023a). Nevertheless, Transformer-based models(Zhou et al. 2021; Wu et al. 2021; Wang et al. 2024a) have emerged as a powerful paradigm by utilizing self-attention mechanisms(Vaswani et al. 2017) to model superior long-term dependency across long sequences, so that effectively capturing persistent temporal relationships. However, Transformer-based models face substantive criticism regarding permutation invariance and quadratic computational complexity limitations in recent researches(Zeng et al. 2023; Tang and Zhang 2024). While significant progress has been made, existing deep architectures primarily operate on single or fixed decomposition, and often model dependencies in a fragmented manner, lacking a mechanism for dynamic, adaptive multi-scale decomposition and coordinated dependency modeling.
@@
To effectively address intricate temporal dependencies, most studies have adopted multiscale modeling perspective, employing diverse multiscale decomposition operations on time series. TimesNet(Wu et al. 2023a) learns frequency-adaptive periods in the frequency domain, transforming time series into 2D tensors to explicitly model intra-period and inter-period variations. TimeMixer(Wang et al. 2024b) decomposes sequences into seasonal and trend components(Wu et al. 2021), subsequently applying different granularity mixing strategies to integrate multiscale features. However, simply using moving average inadequately preserves local semantic patterns. Diverging from the above operations, some methods focus on variable-wise and patch-wise disentanglement. PatchTST(Nie et al. 2023) segments series into subseries-level patches as input tokens, employing channel independence to capture local semantics. ITransformer(Liu et al. 2024) embeds entire series as singular tokens to model extended global representations. TimeXer(Wang et al. 2024d) hierarchically represents variables through dual disentanglement, variable-wise for cross-channel interactions and patch-wise for endogenous variables. However, these decomposition techniques are constrained by static or predefined scale configurations, which hinders their adaptability to complex temporal patterns.
Figure 1: The framework of DMSC. EMPD dynamically segments input series into hierarchical patches, TIB jointly models three dependencies, and ASR-MoE fuses multiscale predictions adaptively.
Multiscale Prediction Operations in TSF
Meanwhile, researches also focuses on multiscale operations in prediction heads. TimeMixer(Wang et al. 2024b) employs dedicated projectors-one per scale-to generate scale-specific predictions. Diverging from simplistic additive aggregation, TimeMixer++(Wang et al. 2025) utilizes frequency-domain amplitude analysis to weightedly aggregate projector outputs. TimeKAN(Huang et al. 2025) adopts frequency transformation and padding operations to unify multiscale representations into identical dimensions, enabling holistic scale integration. Current multiscale fusion mechanisms predominantly rely on static weighting schemes or fixed designs. They fail to dynamically prioritize different scales with intricate temproal dependencies. In contrast to these fragmented approaches, DMSC pioneers fully-spectrum multiscale coordination spanning embedding, extraction, and prediction stages, achieving optimal balance between computational efficiency and forecasting accuracy.
@@
Time series forecasting addresses the fundamental challenge of predicting future values ${Y}={y_{t+1},...,y_{t+\tau}}\in\mathbf{R}^{\tau\times{C}}$ from historical observations ${X}={x_{1},...,x_{t}}\in\mathbf{R}^{t\times{C}}$, where $\tau$ is the prediction horizon, $t$ is the history horizon, and $C$ is the number of variate. To model complex multiscale dependencies inherent in real-world series, the Dynamic Multi-Scale Coordination (DMSC) framework is proposed. As illustrated in Fig. 1, DMSC establishes full-spectrum multi-scale coordination across embedding, feature extraction, and prediction stages using EMPD, TIB and ASR-MoE.
@@
The multi-layer progressive cascade architecture enables hierarchical feature learning through stacked EMPD-TIB units, creating a dynamic feature pyramid where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated pathways. Formally, given input $\mathbf{X}\in\mathbf{R}^{C\times L}$, layer $l$ processes:
@@
where $\mathcal{T}$ is the set of hierarchical features generated by this architecture for ASR-MoE, with each $\mathbf{F}_{l}\in\mathbf{R}^{C\times{D}}$ representing the $l$-th layer’s output. $\mathcal{G}_{l}$ is a gated projection matrix that adaptively modulates the residual flow. This cascade flow enables progressive refinement of multiscale representations, where $\mathbf{F}_{l-1}$ adaptively guides $\mathbf{Z}_{l}$ with residual connections, and $\textbf{TIB}_{l}$ iteratively enhances representations by jointly integrating intra-patch, inter-patch, and cross-variable dependencies. By interleaving dynamic patch decomposition (EMPD) with triadic dependency modeling (TIB), our cascade architecture achieves input-aware progressive refinement of temporal granularities, enabling coherent modeling of both short-term dynamics and long-range trends while maintaining parameter efficiency.
@@
Patch decomposition serves as a fundamental operation to transform time series into structured representations that preserve local semantics while enabling hierarchical pattern discovery. Unlike conventional single-fixed or fixed-scale approaches, the EMPD block introduces input-adaptive and hierarchical decomposition mechanism as a built-in component, dynamically adjusting patch granularities based on sequence characteristics. Thus eliminating rigid predefined patch configurations while natively integrating multiscale decomposition into the model architecture.
Specifically, given an input time series tensor $\mathbf{X}\in\mathbf{R}^{C\times{L}}$(where $C$ is the number of features, $L$ is the sequence length), EMPD first computes a scale factor $\alpha$ through a lightweight neural controller:
@@
where $\mathcal{N}_{\theta}$ is a micro-MLP with sigmoid activation, $\Phi_{\mathrm{GAP}}$ compresses temporal dimensions through averaging. Factor $\alpha\in[0,1]$ dynamically determines the base patch length, and then the patch length decays exponentially as follows:
$\mathbf{P}_{{base}}=[\mathbf{P}_{min}+\alpha\cdot(\mathbf{P}_{max}-\mathbf{P}_{min})],$
$\mathbf{P}_{l}=\mathrm{max}(\mathbf{P}_{min},[\mathbf{P}_{base}/\tau^{l}]),$ (3)
where $\mathbf{P}_{max}$ and $\mathbf{P}_{min}$ denoting the learned parameter for minimum and maximum patch bounds, and $\mathbf{P}_{l}$ means patch length in layer $l$, which ensures shallow layers process coarse-grained dependencies while deeper layers capture fine-grained patterns. Then EMPD applies replication padding to handle boundary effects, followed by hierarchical patch unfolding. The operations are formalized as:
$\mathbf{X}_{p}^{l}=\mathrm{Unfold}(\mathrm{Padding}_{S_{l}}(X)),$ (4)
where $\mathrm{Padding}_{S_{l}}(\cdot)$ denotes replication padding with stride $S_{l}=P_{l}/2$ and $\mathrm{Unfold}_{P_{l},S_{l}}$ unfolds the padded sequence into patches of length $P_{l}$ with stride $S_{l}$. This operation yields a 3D patch tensor $\mathbf{X}_{u}^{l}\in\mathbf{R}^{C\times{N_{l}}\times{P_{l}}}$. Then EMPD transforms $\mathbf{X}_{u}^{l}$ into a unified embedding space via linear projection:
@@
This direct linear projection avoids redundant flattening operations, efficiently transforming patch-wise temporal features into a unified embedding space while preserving the hierarchical structure of multiscale patches.
@@
TIB is designed to model heterogeneous dependencies within the multiscale patch representations generated by EMPD, through joint capture of intra-patch, inter-patch, and cross-variable features. Given the input tensor $\mathbf{Z}^{l}\in{R}^{C\times{N_{l}}\times{D}}$, TIB integrates three complementary dependency types in a coherent framework. Specifically, TIB first processes intra-patch dependencies using depth-wise separable convolutions via $\mathbf{\zeta}_{intra}$ to capture fine-grained local patterns:
$\mathbf{F}_{{intra}}^{l}=\mathbf{\zeta}_{intra}(\mathbf{Z}^{l}),$ (6)
where $\mathbf{\zeta}_{intra}$ consists the depth-wise convolution and $\mathrm{Conv1d}$ to captures local temporal contexts and project features back to the embedding space. This operation preserves intra-patch continuity while extracting localized representations.
@@
$\mathbf{F}_{{inter}}^{l}=\mathbf{\zeta}_{inter}(\mathbf{F}_{{intra}}^{l}),$ (7)
where $\mathbf{\zeta}_{inter}$ consists the dilated convolution and adaptive pooling to capture broader temporal contexts and patch-level information without increasing computational cost.
Then TIB models dependencies between different features by generating adaptive weighting coefficients $\mathbf{G}^{l}\in\mathbf{R}^{{C}\times{1}}$ from globally aggregated inter-patch features $\bar{\mathbf{F}}_{{inter}}^{l}\in\mathbf{R}^{{1}\times{D}}$ as follows:
$\mathbf{G}^{l}=\sigma\left(\mathrm{MLP}(\bar{\mathbf{F}}_{{inter}}^{l})\right),$
$\mathbf{F}_{{cross}}^{l}=\mathbf{G}^{l}\odot\mathbf{F}_{{inter}}^{l},$ (8)
where $\bar{\mathbf{F}}_{{inter}}^{l}$ denotes the global average of $\mathbf{F}_{{inter}}^{l}$, and the sigmoid function $\sigma$ act as a feature-specific gate. The weights $\mathbf{G}^{l}$ are then applied element-wise to $\mathbf{F}_{{inter}}^{l}$ through multiple operation $\odot$, which adaptively scales each feature’s contribution based on its own global relevance, enabling context-aware cross-variable interactions.
Finally, three dependency representations are adaptively fused via learned gating weights, and the fused features are then integrated with residual connection and normalization:
$\mathbf{F}^{l}_{{fused}}=\sum_{i=1}^{3}\mathbf{G}^{l}_{i}\odot\mathbf{F}^{l}_{i},$
$\mathbf{F}^{l}_{{output}}=\mathrm{LayerNorm}(\mathbf{F}^{l}_{{fused}}+\mathbf{Z}^{l}),$ (9)
where $\mathbf{G}^{l}_{i}$ is learned gating weights for each representations. This architecture enables TIB to dynamically balance contributions of different dependency types based on input characteristics, thus effectively capturing intricate dependencies, providing rich temporal features for multiscale prediction.

### v4

@@ Abstract
Time Series Forecasting (TSF) faces persistent challenges in modeling intricate temporal dependencies across different scales. Despite recent advances leveraging different decomposition operations and novel architectures based on CNN, MLP or Transformer, existing methods still struggle with static decomposition strategies, fragmented dependency modeling, and inflexible fusion mechanisms, limiting their ability to model intricate temporal dependencies. To explicitly solve the mentioned three problems respectively, we propose a novel Dynamic Multi-Scale Coordination Framework (DMSC) with Multi-Scale Patch Decomposition block (EMPD), Triad Interaction Block (TIB) and Adaptive Scale Routing MoE block (ASR-MoE). Specifically, EMPD is designed as a built-in component to dynamically segment sequences into hierarchical patches with exponentially scaled granularities, eliminating predefined scale constraints through input-adaptive patch adjustment. TIB then jointly models intra-patch, inter-patch, and cross-variable dependencies within each layer’s decomposed representations. EMPD and TIB are jointly integrated into layers forming a multi-layer progressive cascade architecture, where coarse-grained representations from earlier layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. And ASR-MoE dynamically fuses multi-scale predictions by leveraging specialized global and local experts with temporal-aware weighting. Comprehensive experiments on thirteen real-world benchmarks demonstrate that DMSC consistently maintains state-of-the-art (SOTA) performance and superior computational efficiency for TSF tasks.
@@ Introduction
Time Series Forecasting (TSF) constitutes a pivotal capability across numerous fields such as energy consumption (Alvarez et al. 2010; Guo et al. 2015), healthcare monitoring (Tran, Nguyen, and Shahabi 2019; Wei et al. 2022), transportation scheduling (Guo et al. 2020; Jin et al. 2021), weather forecasting (Bi et al. 2023; Wu et al. 2023b) and economics (Chen et al. 2023b; Yu et al. 2023). Time series data exhibit intricate temporal dependencies, which include nonlinear relationships, multi-scale patterns (e.g., trends and seasonality), and dynamic variable couplings (Qiu et al. 2024). Due to the inherent non-stationarity (Nason 2006) and complex interconnections (Dagum and Bianconcini 2016), traditional approaches often fail to adequately capture the underlying patterns in time series data, making effectively dependency modeling a critical challenge for enhancing predictive accuracy (Shao et al. 2025). In recent years, deep learning has demonstrated a strong capability to capture complex dependency relationships in time series data, making it a formidable tool for TSF (Chen et al. 2023a).
@@
However, intricate dependencies in time series inherently manifest across multiscales: coarser granularities typically encapsulate long-term trends while finer resolutions capture short-term fluctuations and periodic patterns. Conventional single-scale modeling approaches often fail to balance fine-grained local details and holistic global trends (Wang et al. 2025). To effectively model multi-scale temporal dependencies, recent research has pioneered multi-scale modeling frameworks that concurrently extract global and local patterns, thereby thereby mitigating the representational constraints inherent in single-scale approaches. While previous implementations often employ moving average downsampling (Wang et al. 2024b; Huang et al. 2025), which remains insufficient in capturing local semantic information. In contrast, more effective alternatives focus on patch-wise and variable-wise decomposition operations that decompose sequence across feature and temporal dimensions. Such decomposition operations better preserve both global trends and local semantics, and multi-length patches inherently facilitate the learning of multi-scale representations. Critically, prevailing methods suffer from two limitations: 1) reliance on fixed-scale decomposition strategies that lack dynamic scale adaptation; 2) fragmented processing of temporal dependencies and cross-variable interactions, which compromises holistic dependency modeling.
@@
The primary advantage of multi-scale feature extraction lies in its capacity to mine complementary information across diverse temporal granularities and hierarchical levels, thereby overcoming the limitations of single-scale modeling. Following feature capture, the effective design of prediction heads for multi-scale fusion is equally important. Conventional approaches typically employ linear projection layers (Li, Li, and Diao 2025), additive combinations of multiple projections (Wang et al. 2024b), or spectral amplitude-weighted aggregation (Wu et al. 2023a). However, these simplistic fusion methods exhibit fundamental inadequacies: they fail to leverage heterogeneous dominance patterns across multiscales, disregard dynamic inter-scale importance, and incur quadratic growth of model parameters with increasing input lengths and scale counts, thereby severely compromising deployment efficiency (Qiu et al. 2024).
@@
To effectively modeling complex temporal dependencies which inherently manifest across multiple patterns and scales, we propose a Dynamic Multi-Scale Coordination framework (DMSC) for TSF. This framework dynamically decomposes time series and extracts features across scales. It effectively captures intricate dependencies across temporal resolutions and cross-variable interactions. Meanwhile, it can generate adaptive predictions based on learned representations via temporal-aware fusion. Specifically, DMSC processes intricate dependencies through a multi-layer progressive architecture and an Adaptive Scale Routing Mixture-of-Experts (ASR-MoE). The multi-layer progressive cascade architecture incorporates Embedded Multi-Scale Patch Decomposition Block (EMPD), Triad Interaction Block (TIB), in which EMPD transforms the input into 3D representations at varying granularities and TIB models heterogeneous dependencies within each layer. In this architecture, coarse-grained representations from shallow layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. After multi-scale feature extraction, ASR-MoE performs temporal-aware weighted dynamic fusion for multi-scale prediction, thus enabling adaptive integration of cross-scale features to enhancing forecasting accuracy. DMSC formulates a unified framework that incorporates dynamic patch decomposition, deeper models, and sparse MoE principles to achieve full-spectrum multi-scale coordination. This dynamic synergy enables the DMSC framework to achieve SOTA performance across 13 real-world benchmarks while maintaining high efficiency and low computational cost. Our contributions are as follows:
@@
  • A novel multi-layer progressive cascade architecture is designed to jointly integrate Embedded Multi-Scale Patch Decomposition Block (EMPD) and Triad Interaction Block (TIB) across layers. Unlike fixed-scale approaches, EMPD dynamically adjusts patch granularities by employing a lightweight network based on the temporal characteristics of input data, and TIB jointly models intra-patch, inter-patch, and inter-variable dependencies through gated feature fusion to form a coarse-to-fine feature pyramid, where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated residual pathways.
@@
  • Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) is proposed to resolve static fusion limitations in prediction. It establishes a hierarchical expert architecture to explicitly decouples long-term and short-term dependencies, in which global-shared experts capture common long-term temporal dependencies, while local-specialized experts model different short-term variations. Significantly, a temporal-aware weighting aggregator is designed to dynamically compute scale-specific prediction contributions with historical memory.
@@
Existing deep models for TSF tasks can be broadly categorized into MLP-based, CNN-based, and Transformer-based architectures (Wang et al. 2024c). MLP-based models (Chi et al. 2021; Oreshkin et al. 2020; Yi et al. 2023) typically capture temporal dependencies through predefined decomposition and fully-connected layers. While these models demonstrate efficiency in capturing intra-series patterns, they exhibit notable limitations in modeling long-term dependencies and complex inter-series relationships. In contrast, CNN-based models (Bai, Kolter, and Koltun 2018; LIU et al. 2022; Wang et al. 2023) leverage different convolutions to effectively capture local dependencies. Although they remain constrained in modeling global relationships, recent efforts have mitigated this through large-kernel convolutions (Luo and Wang 2024) and frequency-domain analysis (Wu et al. 2023a). Nevertheless, Transformer-based models (Wang et al. 2024a; Wu et al. 2021; Zhou et al. 2021) have emerged as a powerful paradigm, leveraging self-attention mechanisms (Vaswani et al. 2017) to effectively capture long-range dependencies and persistent temporal relationships. However, Transformer-based models face substantial criticism due to their permutation invariance and quadratic computational complexity limitations in recent researches (Tang and Zhang 2024; Zeng et al. 2023). Despite significant progress, most existing deep architectures rely on single or fixed decomposition strategies, and often model dependencies in a fragmented manner, lacking mechanisms for dynamic, adaptive multi-scale decomposition and coordinated dependency modeling.
@@
To effectively capture intricate temporal dependencies, numerous studies have adopted a multi-scale modeling perspective, employing diverse decomposition operations on time series. TimesNet (Wu et al. 2023a) learns frequency-adaptive periods in the frequency domain and transforms series into 2D tensors, explicitly modeling intra-period and inter-period variations. TimeMixer (Wang et al. 2024b) decomposes sequences into seasonal and trend components (Wu et al. 2021), then applies mixing strategies at different granularities to integrate multi-scale features. However, simply using moving average inadequately preserves local semantic patterns. Diverging from these operations, several methods focus on variable-wise and patch-wise disentanglement. PatchTST (Nie et al. 2023) segments series into subseries-level patches as input tokens and employs channel independence to capture local semantics. ITransformer (Liu et al. 2024) embeds entire series as single tokens to model extended global representations. TimeXer (Wang et al. 2024d) hierarchically represents variables through dual disentanglement, variable-wise for cross-channel interactions and patch-wise for endogenous variables. However, these decomposition techniques are constrained by static or predefined scale settings, hindering their adaptability to complex temporal patterns.
Figure 1: The framework of DMSC. EMPD dynamically segments input series into hierarchical patches, TIB jointly models three dependencies, and ASR-MoE fuses multi-scale predictions adaptively.
Multi-Scale Prediction Operations in TSF
Beyond multi-scale decomposition operations, researchers have also investigated multi-scale operations in prediction heads. TimeMixer (Wang et al. 2024b) employs dedicated projectors to generate scale-specific predictions. Diverging from simplistic additive aggregation, TimeMixer++ (Wang et al. 2025) utilizes frequency-domain amplitude analysis to perform weighted aggregation of projector outputs. TimeKAN (Huang et al. 2025) adopts frequency transformation and padding operations to unify multi-scale representations into identical dimensions, thereby enabling holistic scale integration. Current multi-scale fusion mechanisms predominantly rely on static weighting or fixed designs. They fail to dynamically prioritize the importance of different scales with intricate temporal dependencies. Different from these fragmented approaches, our proposed DMSC framework explores full-spectrum multi-scale coordination across embedding, extraction, and prediction stages, achieving both superior forecasting accuracy and high computational efficiency.
@@
Time series forecasting addresses the fundamental challenge of predicting future values ${Y}={y_{t+1},...,y_{t+h}}\in\mathbb{R}^{h\times{C}}$ from historical observations ${X}={x_{1},...,x_{t}}\in\mathbb{R}^{t\times{C}}$, where $h$ is the prediction horizon, $t$ is the history horizon, and $C$ is the number of variables. To model complex multi-scale dependencies inherent in real-world time series, we propose the Dynamic Multi-Scale Coordination (DMSC) framework. As illustrated in Fig. 1, DMSC establishes full-spectrum multi-scale coordination across three key stages: embedding, feature extraction, and prediction, realized through EMPD, TIB and ASR-MoE components.
@@
The multi-layer progressive cascade architecture facilitates hierarchical feature learning through stacked EMPD-TIB units, forming a dynamic feature pyramid. Within this pyramid, coarse-grained representations from earlier layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. Formally, given input $\mathbf{X}\in\mathbb{R}^{C\times L}$, the $l$-th layer processes:
@@
where $\mathcal{T}$ denotes the set of hierarchical features generated by this architecture for ASR-MoE, with each $\mathbf{F}_{l}\in\mathbb{R}^{C\times{D}}$ representing the output of $l$-th layer. Here, $\mathcal{G}_{l}$ represents a gated projection matrix that adaptively modulates the residual information flow. This cascade flow enables progressive refinement of multi-scale representations, where $\mathbf{F}_{l-1}$ adaptively guides the computation of $\mathbf{Z}_{l}$ with residual connections, and $\textbf{TIB}_{l}$ iteratively enhances representations by jointly modeling intra-patch, inter-patch, and cross-variable dependencies. By combining dynamic patch decomposition (EMPD) with triadic dependency modeling (TIB), our cascade architecture achieves input-aware progressive refinement of temporal granularities. This integration enables coherent modeling of both short-term dynamics and long-range trends while maintaining parameter efficiency.
@@
Patch decomposition serves as a fundamental operation that transforms time series into structured representations, preserving local semantics while enabling hierarchical pattern discovery. Unlike conventional single or fixed-scale approaches, the EMPD block introduces adaptive and hierarchical decomposition mechanism as a built-in component, dynamically adjusting patch granularities based on sequence characteristics. This approach eliminates rigid predefined patch configurations while natively integrating multi-scale decomposition into our framework.
Specifically, given an input time series tensor $\mathbf{X}\in\mathbb{R}^{C\times{L}}$(where $C$ denotes the number of variates and $L$ the sequence length), EMPD first computes a scale factor $\alpha$ through a lightweight neural network:
@@
where $\mathcal{N}_{\theta}$ denotes a lightweight MLP with sigmoid activation, and $\Phi_{\mathrm{GAP}}$ denotes a global average pooling operation that compresses the temporal dimension. The factor $\alpha\in[0,1]$ dynamically determines the base patch length, which subsequently undergoes exponential decay across layers as follows:
$\mathbf{P}_{\text{base}}=[\mathbf{P}_{\text{min}}+\alpha\cdot(\mathbf{P}_{\text{max}}-\mathbf{P}_{\text{min}})],$
$\mathbf{P}_{l}=\mathrm{max}(\mathbf{P}_{\text{min}},[\mathbf{P}_{\text{base}}/\tau^{l}]),$ (3)
where $\mathbf{P}_{\text{max}}$ and $\mathbf{P}_{\text{min}}$ denote the parameters for minimum and maximum patch bounds, respectively. $\mathbf{P}_{l}$ represents the patch length at the $l$-th layer. This design ensures that shallow layers process coarse-grained dependencies while deeper layers capture fine-grained ones. EMPD then applies replication padding to mitigate boundary effects, followed by hierarchical patch unfolding operation:
$\mathbf{X}_{p}^{l}=\mathrm{Unfold}_{P_{l},S_{l}}(\mathrm{Padding}_{S_{l}}(X)),$ (4)
where $\mathrm{Padding}_{S_{l}}(\cdot)$ denotes the replication padding operation with stride $S_{l}=P_{l}/2$ and $\mathrm{Unfold}_{P_{l},S_{l}}$ unfolds the padded sequence into patches of length $P_{l}$ using a stride of $S_{l}$. This operation produces a 3D patch tensor $\mathbf{X}_{u}^{l}\in\mathbb{R}^{C\times{N_{l}}\times{P_{l}}}$. Finally, EMPD projects $\mathbf{X}_{p}^{l}$ into a unified embedding space via a linear projection:
@@
This direct linear projection efficiently maps patch-wise temporal features into a unified embedding space, avoiding redundant flattening operations while preserving the hierarchical structure of the multi-scale patches.
@@
The Triad Interaction Block(TIB) is designed to model heterogeneous dependencies within the multi-scale patch representations generated by EMPD, through joint capture of intra-patch, inter-patch, and cross-variable interactions. Given the input tensor $\mathbf{Z}^{l}\in\mathbb{R}^{C\times{N_{l}}\times{D}}$, TIB integrates these three complementary dependency types within a coherent framework. Specifically, TIB first processes intra-patch dependencies using depth-wise separable convolutions through operation $\mathbf{\zeta}_{\text{intra}}$ to capture fine-grained local patterns:
$\mathbf{F}_{\text{intra}}^{l}=\mathbf{\zeta}_{\text{intra}}(\mathbf{Z}^{l}),$ (6)
where $\mathbf{\zeta}_{\text{intra}}$ comprises depth-wise convolution and $\mathrm{Conv1d}$ to capture local temporal contexts and project features back to the embedding space. This operation preserves intra-patch continuity while effectively extracting local representations.
@@
$\mathbf{F}_{\text{inter}}^{l}=\mathbf{\zeta}_{\text{inter}}(\mathbf{F}_{\text{intra}}^{l}),$ (7)
where $\mathbf{\zeta}_{\text{inter}}$ consists of the dilated convolution and adaptive pooling operations, capturing broader temporal contexts and patch-level information without increasing computational cost.
TIB then models cross-variable dependencies by generating adaptive weighting coefficients $\mathbf{G}^{l}\in\mathbb{R}^{{C}\times{1}}$ as follows:
$\mathbf{G}^{l}=\sigma\left(\mathrm{MLP}(\bar{\mathbf{F}}_{\text{inter}}^{l})\right),$
$\mathbf{F}_{\text{cross}}^{l}=\mathbf{G}^{l}\odot\mathbf{F}_{\text{inter}}^{l},$ (8)
where $\bar{\mathbf{F}}_{\text{inter}}^{l}$ denotes the global average of $\mathbf{F}_{\text{inter}}^{l}$, and the sigmoid function $\sigma$ acts as a feature-specific gate mechanism. $\mathbf{G}^{l}$ are applied element-wise to $\mathbf{F}_{\text{inter}}^{l}$ via the Hadamard product $\odot$, adaptively scaling each feature’s contribution based on its global relevance and enabling context-aware cross-variable interactions.
Finally, the three dependency representations are adaptively fused using learned gating weights, and the resulting features are integrated with residual connection and normalization:
$\mathbf{F}^{l}_{\text{fused}}=\sum_{i=1}^{3}\mathbf{G}^{l}_{i}\odot\mathbf{F}^{l}_{i},$
$\mathbf{F}^{l}_{\text{output}}=\mathrm{LayerNorm}(\mathbf{F}^{l}_{\text{fused}}+\mathbf{Z}^{l}),$ (9)
where $\mathbf{G}^{l}_{i}$ represents the learned gating weights for the $i$-th representations. This architecture enables TIB to dynamically balance the contributions of different dependency types based on the temporal characteristics of input data. Consequently, it effectively captures intricate dependencies and provides rich temporal features for multi-scale prediction.